4. You have prepared a batch of competent E coli cells and found its transformation efficiency is 106 per μg plasmid DNA. A professional scientists prepared competent E. coli cells having a transformation efficiency of about 108 cells/ g plasmid DNA Calculate % score of your E. coli transformation efficiency (Τ.Ε.).
Answer #1

The % of Transformation Efficiency = Practical yield / Theorical yield * 100.

The practical yield = 10^6/ ug and
the theoritical yield = 10^8/ ug.

So, The TE = 10^6/10^8 * 100 = 1.

so, the answer is 1.